OpenStudy (anonymous):

please help!! 1 medal and fan Which is the equation of the line that contains points (8, 10) and (–4, 2)? y-10=2/3(x-8) y-8=3/2(x-10) y-10=3/2(x-8) y-8=2/3(x-10)

OpenStudy (goldeverything):

the last one

OpenStudy (anonymous):

can you please explain your answer

OpenStudy (itsonlycdeee):

The slope of a line is m = y2 - y1 / x2 - x1. Therefore 2-10/ -4-8 = -8/-12 = 2/3. Which slope matches that out of the 4 equations you posted?

OpenStudy (anonymous):

the last one

OpenStudy (anonymous):

first one

OpenStudy (itsonlycdeee):

It's actually the first one because after finding the slope, you use one set of the points they gave you.

OpenStudy (anonymous):

thank you

OpenStudy (itsonlycdeee):

No problem!